# Probing Einstein-Maxwell-Scalar Black hole via Thin Accretion Disks and Shadows with EHT Observations of M87* and Sgr A*

### v2

@@ IV Conclusions
On the other hand, we analyzed the photon trajectories of the Schwarzschild BH, RN BH, and EMS BH. The differences among these three types of BHs are only reflected in the critical impact parameter $b_{c}$ (the BH shadow radius). We found that the Schwarzschild BH has the largest shadow radius, while the EMS BH has the smallest. All three types of BHs undergo a transition from direct emission to the photon ring, then to the lensing ring, and finally back to direct emission as the impact parameter $b$ increases. A similar effect has been observed using the Okyayvg Mathematical Notebook package 182 in our investigation of BH shadows and rings with three toy models of thin accretion disk. Our analysis reveals that the observed intensity is primarily dominated by direct emission, while the lensing ring contributes only marginally to the total flux, and the photon ring remains negligible in all cases. Among the three types of BHs, the Schwarzschild BH exhibits the highest emission intensity, followed by the RN BH, with the EMS BH having the lowest intensity. In other words, an increase in the model parameters leads to a reduction in the disk’s radiation intensity.

### v3

@@ IV Conclusions
On the other hand, we analyzed the photon trajectories of the Schwarzschild BH, RN BH, and EMS BH. The differences among these three types of BHs are only reflected in the critical impact parameter $b_{c}$ (the BH shadow radius). We found that the Schwarzschild BH has the largest shadow radius, while the EMS BH has the smallest. All three types of BHs undergo a transition from direct emission to the photon ring, then to the lensing ring, and finally back to direct emission as the impact parameter $b$ increases. A similar effect has been observed using the Okyayvg Mathematical Notebook package 182 in our investigation of BH shadows and rings with three toy models of thin accretion disk. Our analysis reveals that the observed intensity is primarily dominated by direct emission, while the lensing ring contributes only marginally to the total flux, and the photon ring remains negligible in all cases. Among the three types of BHs, the Schwarzschild BH exhibits the highest emission intensity, followed by the RN BH, with the EMS BH having the lowest intensity. In other words, an increase in the model parameters leads to a reduction in the disk’s radiation intensity. Fianlly, when we consider weak gravitational lensing in EMS BH, our results indicate that for fixed $q$, the total magnification grows with $\alpha$ but decreases with $\beta$ and with increasing BH charge. Moreover, $\beta$ shifts the curves downward when $\alpha=0$, while for $\beta=0$ the trend reverses.
@@ Acknowledgements.
Appendix A Derivations for $r_{ph}$
For a general static, spherically symmetric spacetime, the photon sphere radius $r_{\mathrm{ph}}$ can be determined by the extremum condition of the effective potential for null geodesics. Using the metric
$ds^{2}=-f(r)dt^{2}+f^{-1}(r)dr^{2}+C(r)\left(d\theta^{2}+\sin^{2}\theta\,d\phi^{2}\right),$ (49)
the condition for the photon sphere reads
$\frac{d}{dr}\left(\frac{C(r)}{f(r)}\right)=0,$ (50)
or equivalently:
$f^{\prime}(r)\,C(r)-f(r)\,C^{\prime}(r)=0.$ (51)
In our EMS black hole model, the metric functions are
$\begin{cases}f(r)=\left(1-\frac{b_{1}}{r}\right)\left(1-\frac{b_{2}}{r}\right)^{\frac{1-\alpha^{2}}{1+\alpha^{2}}}+\frac{\beta Q^{2}}{C(r)},\\ C(r)=r^{2}\left(1-\frac{b_{2}}{r}\right)^{\frac{2\alpha^{2}}{1+\alpha^{2}}},\end{cases}$ (52)
with
$\begin{cases}b_{1}=\left(1+\sqrt{1-q^{2}(1-\alpha^{2})}\right)M,\quad\\ b_{2}=\frac{1+\alpha^{2}}{1-\alpha^{2}}\left[1-\sqrt{1-q^{2}(1-\alpha^{2})}\right]M.\end{cases}$ (53)
Now, for simplicity, we define
$p=\frac{1-\alpha^{2}}{1+\alpha^{2}},\quad s=\frac{2\alpha^{2}}{1+\alpha^{2}},$ (54)
then we have
$\begin{cases}C^{\prime}(r)&=2r\left(1-\frac{b_{2}}{r}\right)^{s}+sb_{2}\left(1-\frac{b_{2}}{r}\right)^{s-1},\\ f^{\prime}(r)&=\frac{b_{1}}{r^{2}}\left(1-\frac{b_{2}}{r}\right)^{p}+\left(1-\frac{b_{1}}{r}\right)p\left(1-\frac{b_{2}}{r}\right)^{p-1}\frac{b_{2}}{r^{2}}-\beta Q^{2}\frac{C^{\prime}(r)}{C(r)^{2}}.\end{cases}$ (55)
Substituting $f(r),f^{\prime}(r),C(r)$ and $C^{\prime}(r)$ into the photon sphere condition
$f^{\prime}(r)\,C(r)-f(r)\,C^{\prime}(r)=0,$ (56)
we obtain the explicit nonlinear equation for $r_{\mathrm{ph}}$ in terms of the model parameters $\alpha,\beta,Q,M$,
$0=$ ${\left[\frac{b_{1}}{r^{2}}\left(1-\frac{b_{2}}{r}\right)^{p}+\left(1-\frac{b_{1}}{r}\right)p\left(1-\frac{b_{2}}{r}\right)^{p-1}\frac{b_{2}}{r^{2}}-\beta Q^{2}\frac{2r\left(1-\frac{b_{2}}{r}\right)^{s}+sb_{2}\left(1-\frac{b_{2}}{r}\right)^{s-1}}{r^{4}\left(1-\frac{b_{2}}{r}\right)^{2s}}\right]r^{2}\left(1-\frac{b_{2}}{r}\right)^{s}}$ (57)
$-\left[\left(1-\frac{b_{1}}{r}\right)\left(1-\frac{b_{2}}{r}\right)^{p}+\frac{\beta Q^{2}}{r^{2}\left(1-\frac{b_{2}}{r}\right)^{s}}\right]\times\left[2r\left(1-\frac{b_{2}}{r}\right)^{s}+sb_{2}\left(1-\frac{b_{2}}{r}\right)^{s-1}\right].$
This equation is highly nonlinear and cannot be solved analytically in closed form for arbitrary parameter values. Therefore, the photon sphere radius can only be expressed implicitly.
However, once specific values of $\alpha,\beta,Q,M$ are chosen, the implicit equation can be solved numerically to yield $r_{\mathrm{ph}}$ with high precision using standard computational tools. This approach allows for direct investigation of the dependence of the photon sphere radius on the physical and coupling parameters of the EMS black hole.
Appendix B Derivations for $\hat{\alpha}$
We follow the Asahi Ishihara’s method 666 to calculate the deflection angle. At first, we have
$dt^{2}=\gamma_{ij}dx^{i}dx^{j},$ (58)
where $i$ and $j$ denote 1,2 and 3 , and $\gamma_{ij}$ is often called the optical metric. Then we can have the Gaussian Curvature
$K=\frac{R_{r\phi r\phi}}{g},\quad g=\det(\gamma_{ij})=\frac{C(r)}{f(r)^{3}},$ (59)
and the Gaussian Curvature then can be written as
$K=-\frac{1}{2}\sqrt{\frac{f^{3}(r)}{C(r)}}\frac{\partial}{\partial r}\left(\sqrt{\frac{f^{3}(r)}{C(r)}}\frac{\partial}{\partial r}\left(\frac{C(r)}{f(r)}\right)\right),$ (60)
thus (in the weak limit $M\ll r$)
$K\approx$ $\frac{1}{r^{6}\left(r^{2}-Mrq^{2}\alpha^{2}\right)^{2}}\left(M(3M-2r)r^{6}+Mr^{2}rq^{2}\left(r^{4}+3M^{3}rq^{2}-2M^{2}r\left(r+rq^{2}\right)\right)\alpha^{2}\right$. (61)
$+Q^{2}(2M-r)r\left(3r^{4}-4Mr^{2}rq^{2}\alpha^{2}+3M^{2}rq^{4}\alpha^{4}\right)\left(-1+\alpha^{2}-\beta\right)$
$\left.+Q^{4}\left(2r^{4}-3Mr^{2}rq^{2}\alpha^{2}+2M^{2}rq^{4}\alpha^{4}\right)\left(1-\alpha^{2}+\beta\right)^{2}\right).$
Next, we will use these equations to find the weak deflection angle using the Gauss-Bonnet theorem (GBT), originally stated as 667 ; 668
$\iint_{D}KdS+\sum_{a=1}^{N}\int_{\partial D_{a}}\kappa_{\mathrm{g}}d\ell+\sum_{a=1}^{N}\theta_{a}=2\pi\chi(D).$ (62)
[…]
It was shown by 666 that in a SSS spacetime admitting asymptotic flatness, the deflection angle can be written as
$\hat{\alpha}=-\iint_{D}K\,dS.$ (63)
Taking the half-plane $\phi\in[0,\pi]$ with area element
$dS\approx r+M\left(3-\frac{a^{2}q^{2}}{2}\right)+\frac{1}{r}\left[-\frac{3}{2}Q^{2}\left(1-a^{2}+\beta\right)+M^{2}\left(\frac{15}{2}-\frac{3a^{2}q^{2}}{2}-\frac{a^{4}q^{4}}{8}\right)\right]dr\,d\phi,$ (64)
integrate over $r\in[r_{0},\infty]$.
The final result is
$\hat{\alpha}\approx\frac{4M}{b}-\frac{3\pi Q^{2}\left(1-\alpha^{2}+\beta\right)}{4b^{2}}.$ (65)